How to construct a pie chart
NASA Budget Use a pie chart to display the data. The data represents the 2007 NASA budget (in millions of dollars)divided amon three categories.
Science, aeronautics and exploration 10,651
Explortaion 6108
Inspector General 34
Need to know how to convert the numbers (in millions of dollars)to a percent to be able to start the pie chart.

Answer by nerdybill(7384)

How to construct a pie chart
NASA Budget Use a pie chart to display the data. The data represents the 2007 NASA budget (in millions of dollars)divided amon three categories.
Science, aeronautics and exploration 10,651
Explortaion 6108
Inspector General 34
Need to know how to convert the numbers (in millions of dollars)to a percent to be able to start the pie chart.
.
First, find the total:
10651 + 6108 + 34 = 16793
.
Percentages are then:
Science, aeronautics and exploration: 10651/16793 * 100 = 63.43%
Explortaion: 6108/16793 * 100 = 36.37%
Inspector General: 34/16793 * 100 = 0.20%